The expression below can be rewritten in several ways.
(4x-5y)(16x^2+kxy+25y^2)
If this expression can be represented by (a+b)^3 where a and b are algebraic terms, what is the value of k?
A. -40
B. -20
C. 9
D. 40

Answer by ankor@dixie-net.com(22740)

The expression below can be rewritten in several ways.
(4x-5y)(16x^2+kxy+25y^2)
If this expression can be represented by (a+b)^3 where a and b are algebraic terms, what is the value of k?
:
therefore we know
(4x - 5)(4x - 5y) = 16x^2 + kxy + 25y^2
FOIL the left side
16x^2 - 20x - 20x + 25x^2 = 16x^2 + kxy + 25x^2
combine like terms
16x^2 - 16x^2 - 40xy + 25x^2 - 25x^2 = kxy
-40xy = kxy
Divide both sides by xy
-40 = k